The electric potential immediately outside another charged conducting sphere is 240 V, and 10.0 cm above the surface of the sphere the magnitude of the electric field is 390 V/m.

(c) Determine all possible values for the radius of the sphere. (Enter your answers from smallest to largest. If only one value exists, enter "NONE" in the second answer blank.)

r1 = cm
r2 = cm


(d) Determine the charge on the sphere for each value of r. (If only one value exists, enter "NONE" in the second answer blank.)

q1 = nC
q2 = nC
Answer #1

electric potential at the surface of the sphere is

                       V1 = kq/ R  

                    240 = kq/ R

                       kq = 240*R ................... (1)

electric field at (R + 10) from the center of the sphere is

         E = kq/(R + 10)2

     390 = kq/(R + 10)2

       kq = 390(R + 10)2 ..................... (2)

compare eq (1) and (2), we get

        240*R = 390(R + 10)2

        240*R = 390(R + 0.1 m)2

         390 R2 - 162 R + 3.9 = 0

solve the above quadratic equation , we get

      R = 0.3897 m or R = 0.02566 m

       R ˜ 38.97 cm or R ˜ 2.56 cm

radius of the sphere is,

          R = 38.97 cm (or) R = 2.56 cm
if R = 38.97 cm ,
substitute the value R = 38.97 cm in eq (1), we get
charge on the sphere q = 9.38 nC
if R = 2.56 cm ,
substitute the value R = 2.56 cm in eq (1), we get
charge on the sphere q = 0.614 nC